# Smaller is Better: Fusing Entity Embeddings with Low-dimensional Projections of LLM-based Representations

### v1

@@ 5 Conclusions and Further Work
In this work, we propose new knowledge-enriched, LLM-based low-dimensional document embeddings. The results suggest that fusing modalities in low dimensions not only preserves space but also enables efficient representations that surpass even proprietary embeddings. This is in line to an extent with the work of [8] where it was shown that contextual embeddings can be approximated by low-dimensional geometry. We advance the related line of work by introducing sophisticated named entity linking and AutoML while leveraging representations extracted from LLMs. Our findings demonstrate that these embeddings perform well across different datasets and domains, showing promising potential for future applications. The results indicate that even in unsupervised applications, such as clustering, these lightweight embeddings might provide robust document representations. Furthermore, by applying AutoML, we show that learning in low dimensions is feasible and competitive with high-dimensional embeddings.
The implications of these results are that: A1. KG-enriched representations can outperform text representations, including both encoder and decoder-based models (A3), and that learning on these representations in low dimensions is feasible (A2) even for proprietary document representations (A4).
In this study, we utilized only a portion of the BabelNet graph and the WikiData5m subgraph. In the future, we aim to include the entire graph, improve the fusion process with advanced disambiguation methods, and explore injection of external knowledge at the token level to create a synergy between LLMs and KGs, as suggested by Pan [23]. Finally, we plan to explore if recursive compression of our proposed representations provides better down-stream results.
@@ Acknowledgments
The authors acknowledge financial support from the Slovenian Research and Innovation Agency through research core funding (No. P2-0103) and projects No. J4-4555, J5-3102, L2-50070, and PR-12394.

### v2

@@ 5 Conclusions and Further Work
In this work, we propose new knowledge-enriched, LLM-based low-dimensional document embeddings. The results suggest that fusing modalities in low dimensions not only preserves space but also enables efficient representations that surpass even proprietary embeddings. This is in line to an extent with the work of [8] where it was shown that contextual embeddings can be approximated by low-dimensional geometry. We advance the related line of work by introducing sophisticated named entity linking and AutoML while leveraging representations extracted from LLMs. Our findings demonstrate that these embeddings perform well across different datasets and domains, showing promising potential for future applications. The results indicate that, even in unsupervised applications, such as clustering, these lightweight embeddings might provide robust document representations. Furthermore, by applying AutoML, we show that learning in low dimensions is feasible and competitive with high-dimensional embeddings.
The implications of these results are that: A1. KG-enriched representations can outperform text representations, including both encoder and decoder-based models (A3), and that learning on these representations in low dimensions is feasible (A2), even for proprietary document representations (A4).
In this study, we utilized only a portion of the BabelNet graph and the WikiData5m subgraph. In the future, we aim to include the entire graph, improve the fusion process with advanced disambiguation methods, and explore injection of external knowledge at the token level to create a synergy between LLMs and KGs, as suggested by Pan et al. [23]. Finally, we plan to explore if recursive compression of our proposed representations provides better down-stream results.
@@ Acknowledgments
The authors acknowledge financial support from the Slovenian Research and Innovation Agency through research core funding (No. P2-0103) and projects No. J4-4555, J5-3102, L2-50070, and PR-12394, and from CREATIVE project (CRoss-modal understanding and gEnerATIon of Visual and tExtual content) funded by the MUR Progetti di Ricerca di Rilevante Interesse Nazionale programme (PRIN 2020).